OpenStudy (anonymous):

Solve for x: 2 over 5 (x - 2) = 4x.

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 } (x-2)=4x\] Is that what your problem looks like?

OpenStudy (anonymous):

yes siree

OpenStudy (anonymous):

@faith2012

OpenStudy (anonymous):

First you would distribute the 2/5 to the x and -2

OpenStudy (anonymous):

It would look like \[\frac{ 2 }{ 5 }x-\frac{ 4 }{ 5 }\]\[=4x\]

OpenStudy (anonymous):

You need to get the x's on one side.

OpenStudy (anonymous):